OpenStudy (anonymous):

3y-5(y-1)=-5 how can I solve this

OpenStudy (anonymous):

what are your thoughts about how it can be solved?

OpenStudy (anonymous):

Well at first I thought solve the one in the perentheseis but I keep getting it wron

OpenStudy (anonymous):

your onthe right track..

OpenStudy (anonymous):

Try distributing the 5 through to each term in the parens, then combine like terms and move the constants to one side and the variable term to the other.

OpenStudy (anonymous):

lets distribute that -5 thru the paranthesis like this: 3y -5y +5 = -5 now what can we do?

OpenStudy (anonymous):

So do I subtract the 3y with the 5y

OpenStudy (anonymous):

you do... 3y -5y = -2y right? -2y +5 = -5 this looking like something you can solve yet?

OpenStudy (anonymous):

i keep getting 6.

OpenStudy (anonymous):

6 as an answer to the problem?

OpenStudy (anonymous):

-2y +5 = -5....what can we do next?

OpenStudy (anonymous):

Thats where I am stuck at. The whole = part confuses me.

OpenStudy (anonymous):

with a basic equation such as this, they want you to become strong in your basic properties of addintion and multiplication. these foundations are what we use in all the other higher orders of math. if you dont have a firm grasp of it now, its gonna hurt you in the future :)

OpenStudy (anonymous):

Yes, I agree. Its hurting me now. I am a college student and its extremely hard

OpenStudy (anonymous):

lets see if we can do the simple step here and I will see if I can explain what we did with that -5 later ok? -2y +5 = -5 we have a balanced equation here. whatever we add to one side has to be added to the other to keep it equal, balanced

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Remember that because each side equals the other, you can maintain equality by doing the same thing on both sides. If we want to isolate (get by itself) the term with the y, we can subtract 5 from both sides.

OpenStudy (anonymous):

what happens when we add (-5) to both sides? -2y +5-5 = -5-5 -2y +0 = -10 -2y = 10 you with me so far?

OpenStudy (anonymous):

-2y = -10. I missed that negative on there sorry.

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

now to get y all alone, we need to divide it by (-2) why? becasue -2/-2 = 1 -2y/-2 = -10/-2 1y = 5 y = 5

OpenStudy (anonymous):

Oh. I see. Thanks a mill

OpenStudy (anonymous):

noprob :)

OpenStudy (anonymous):

Thats was easier than I thought. I was looking too much into it.

OpenStudy (anonymous):

alot of people do, they get nervous with it and what really happens with math is you know HOW to do it, but are afraid to be wrong in the end :) at least thats how I feel alot

OpenStudy (anonymous):

Thats right. Well thanks you I aprreciate your help.